I am not sure if I did this right? If not, please explain how to do it?
(3t-4)(t^-5t+1)
3t*t^2=3t^2
3t*-5t=-15t
3t*1=3t
-4*t^2=-4t^2
-4*-5t=20t
-4*1=-4
3t^2-15t-3t--4t^2-20t--4
Now what? Is this the answer?

Answer by Alan3354(69443)

(3t-4)(t^-5t+1)
3t*t^2=3t^2 ****** It's 3t^3, not t^2
3t*-5t=-15t **** It's -15t^2
3t*1=3t
-4*t^2=-4t^2
-4*-5t=20t
-4*1=-4
3t^2-15t-3t--4t^2-20t--4
Now what? Is this the answer?